What is the SIPB?

The Student Information Processing Board (pronounced ‘Sip´ -bē’, sometimes written as ‘ΣΠB’) is a student organization that supports computing at MIT. Our members are volunteers who range from hard-core hackers to casual users. What we share is a desire to help other students and members of the MIT community with their computing needs.

How can I contact the SIPB?

The SIPB office is on the fifth floor of the Student Center, next to the Athena cluster. Our room number is W20-557, and our phone number is x3-7788. We’re staffed completely by volunteers, many of which are students. This means that we’re open almost any time of the day, including 4 a.m.

What does the SIPB offer?

The SIPB supports all sorts of neat stuff. Our mandate to “support computing at MIT” is fulfilled in many diverse ways, and if you have a new idea, we’d like to hear it. Typically, we answer questions over the phone, on-line, and in person; provide certain resources (such as documentation and access to unusual hardware) to the MIT community; and support hardware and software that “falls between the cracks” of the other support mechanisms at MIT.

In particular, we provide support for the Athena user environment, make available additional software and hardware (such as a digital scanner, various tape drives, and a Macintosh on MITnet). We also provide support for both Linux and NetBSD, working to help make it possible for people to use ‘Athena’ on their home computers. We also provide many services to MITnet and the Internet, including running a news server (news.mit.edu), web server (http://www.mit.edu/), and an Internet FAQ Archive ftp/mail server (rtfm.mit.edu).

Who can join the SIPB?

Anyone who is interested in what we do can become a prospective member of the SIPB. You don’t have to know a lot about computers, and you don’t have to be a computer science major. You do have to want to learn and to help other users. Our meetings are held every Monday night at 7:30 in our office (W20-557), and are open to the public.

If you’re wondering what sort of “help” you’d be giving, just look at the list of things we do. If you can answer questions on the phone or in person, or write documentation or programs, or are interested in learning more about these things, then come by and hang out with us, or take a look at our web page: http://www.mit.edu/sipb/